30% of the population of particular country favour the political party BNM. Suppose that a sample of 15 persons were randomly selected and interviewed. What is the probability that
1) exactly 2 of them would favour BNM?
2)less than 14 of them would favour BNM?
3) What is the variance of the random variable X, where X is the number of persons in the sample who supported BNM?
Suppose a random sample of 120 persons are selected. How many persons should one expect to not be supporters of BNM?

Answer by ewatrrr(24785)


Hi
p = .30, n = 15
1)P(x=2) = 15C2(.30)^2(.8)^13 0r binompdf(15, .30, 2) = .0916 0r 9.16%
2) P(x < 14) = 1 - P(x ≥ 14 = P(x ≤ 13) = .9999 0r 99.99%
3) Variance = (.30x)(.70)
Sample 120
.70(120) = non-supporters = 84